A Norman window consists of rectangle surmounted by a semicircle. If the


perimeter of a Norman window is 32 ft, what should be the radius of the


semicircle and the height of the rectangle such that the window will admit most


light?

1
Expert's answer
2021-11-23T17:13:39-0500

Let x denote half the width of the rectangle (so x is the radius of the semicircle), and let h denote the height of the rectangle.

perimeter of the window:

"2y + 2x +2\u03c0x\/2 = 2y + (2 + \u03c0)x=32"

then:

"y=16-\frac{x(\pi +2)}{2}"


 area of the window:

"A=2xy+\pi x^2/2"

"A(x)=2x(16-\frac{x(\pi +2)}{2})+\pi x^2/2=32x-\frac{x^2(\pi +4)}{2}"

"A'(x)=32-(4+\pi)x=0"


radius of the semicircle such that the window will admit most light:


"x=\frac{32}{4+\pi}=4.48" ft


height of the rectangle such that the window will admit most light:


"y=16-\frac{4.48(\pi +2)}{2}=4.48" ft